# Emotional EEG Classification using Upscaled Connectivity Matrices

### v2

@@ I Introduction
Emotions play a vital role in shaping the user experience in modern intelligent systems, including content recommendation, virtual reality, and human-computer interaction (HCI). Consequently, there is an increasing demand for accurate and real-time emotion recognition methods. Traditional approaches based on self-reports, such as surveys and interviews, suffer from subjectivity and poor temporal resolution. In contrast, electroencephalography (EEG)-based techniques offer real-time monitoring capabilities and reduced bias, making them attractive for implicit emotion analysis [1].
Early EEG-based methods mainly extracted features in the amplitude or frequency domain. Although effective to some extent, these approaches often fail to capture the dynamic interactions between brain regions. To overcome this limitation, brain connectivity analysis has gained attention, focusing on the relationships among different brain regions [2], [3], [4]. Connectivity matrices, constructed using metrics such as Pearson’s correlation coefficient (PCC), phase locking value (PLV) and transfer entropy (TE), provide a two-dimensional representation of interregional interactions and are widely used in emotional EEG classification [5].
@@ II-A EEG-Based Emotion Recognition
EEG has been widely utilized in emotion recognition due to its ability to capture real-time brain activity [1]. Both time-domain and frequency-domain analyses have been employed to interpret EEG signals associated with emotional states.
In particular, frequency-domain features have shown strong relevance to emotion classification. Liu et al. [6] distinguished emotional categories such as joy, anger, fear, and sadness using power spectral density (PSD) and asymmetry indices. Zhang et al. [7] further enhanced classification performance by incorporating time-frequency distribution analysis, which captures both temporal and spectral characteristics of EEG signals.
More recently, advances in deep learning have significantly improved feature extraction from EEG. Among various architectures, CNNs have demonstrated high performance in automated emotion recognition. Yanagimoto and Sugimoto [8] applied CNNs to classify emotional valence and arousal levels, highlighting the effectiveness of learned representations over hand-crafted features. Li et al. [9] proposed transforming EEG scalp maps into 2D images for CNN processing, enabling better handling of spatial and individual variability in emotional responses.
@@ II-B Brain Connectivity in Emotion Analysis
Costa et al. [10] demonstrated that phase synchronization patterns in EEG are significantly correlated with emotional states. Lee and Hsieh [11] proposed a method that integrates correlation, coherence, and phase synchronization to enhance emotion classification accuracy.
The importance of connectivity-based analysis is further supported by recent neuroimaging studies. Saarimäki et al. [12] showed that functional connectivity patterns from fMRI data reliably differentiate discrete emotion categories, especially within the default mode network. In the context of EEG, a recent systematic review by Khaleghi et al. [13] comprehensively analyzed how various brain abnormalities alter functional connectivity patterns, providing a broader clinical perspective on network disruptions.
Building on these findings, recent works have explored the integration of brain connectivity and deep learning. Moon et al. [5] demonstrated that EEG-derived connectivity matrices can serve as effective inputs to CNNs for learning emotion-related features, underscoring the potential of combining neural connectivity representations with data-driven models.
@@ III-A Connectivity Matrix Construction
$\text{PCC}(i,k)=\frac{\frac{1}{T}\sum_{t=1}^{T}(x_{i}^{t}-\mu_{i})(x_{k}^{t}-\mu_{k})}{\sigma_{i}\sigma_{k}}$ (1)1( 1 )
@@
Electrode placement within the connectivity matrix follows the distance-restricted arrangement method proposed in [5], which emphasizes spatial proximity among electrodes to reduce noise from long-range, spurious connections. In other words, electrodes are ordered such that those located in the same hemisphere and in close spatial proximity are positioned adjacently in the matrix. The resulting matrices are processed by CNNs to extract discriminative features for emotion classification.
@@ III-B Limitations of Conventional Approaches
This limitation arises primarily from the intrinsic nature of connectivity matrices, which encode pairwise relationships rather than spatially continuous features. For comparison, Fig. 2 presents a natural image from the CIFAR-10 dataset [14] and its corresponding activation after the first convolutional layer of a CNN trained for object recognition. In this case, the model successfully captures meaningful structures such as edges and textures, as the input contains coherent spatial information. However, in the case of connectivity matrices, the absence of such spatial continuity impedes the CNN’s ability to learn effective representations.
@@ III-C Proposed Method
$r$” refers to the upscaling factor.
@@ IV Experimental Setup
The evaluation in this study utilizes the DEAP dataset [15], which contains EEG signals recorded from 32 participants while they viewed 40 emotional music videos. The EEG signals were initially recorded at a sampling rate of 512 Hz and subsequently downsampled to 128 Hz. The dataset includes emotion labels categorized by valence and arousal scales for each video stimulus. Based on the evaluation framework in [5], we develop a classification model to identify which of the 40 videos was being viewed based on the recorded EEG signals.
@@
As in [5], the one-minute EEG signal for each trial is divided into three-second-long segments. A simple non-overlapping segmentation would yield only 20 segments per trial. To augment the dataset, we apply an overlapping segmentation approach with a step size of 0.5 seconds, thereby increasing the number of samples. For each three-second segment, a 32×32×10 connectivity matrix is computed, capturing inter-channel relationships across the ten sub-frequency bands.
@@ V-B Visualization of t-SNE
To assess the discriminative quality of features extracted by the trained CNNs, we apply t-distributed stochastic neighbor embedding (t-SNE) [16] to the final feature vectors produced using different convolutional kernel sizes ($k=3,5,7$), under a fixed upscaling factor of ×3.0 with nearest-neighbor interpolation. The resulting embeddings are illustrated in Fig. 6.
@@ V-C Visualization of CNN Activation
To further understand how the input resolution affects feature learning, we visualize the activation maps from the first convolutional layer along with their corresponding Grad-CAM [17] outputs, which highlight the regions contributing most to the model’s final prediction. As illustrated in Fig. 6, higher upscaling factors lead to more spatially coherent and sharply defined activation patterns. This observation suggests that increased input resolution allows the model to more effectively exploit the spatial structure inherent in the connectivity matrices.

### v3

@@ I Introduction
Emotions play a vital role in shaping the user experience in modern intelligent systems, including content recommendation, virtual reality, and human-computer interaction (HCI). Consequently, there is an increasing demand for accurate and real-time emotion recognition methods. Traditional approaches based on self-reports, such as surveys and interviews, suffer from subjectivity and poor temporal resolution. In contrast, electroencephalography (EEG)-based techniques offer real-time monitoring capabilities and reduced bias, making them attractive for implicit emotion analysis [moon2016implicit].
Early EEG-based methods mainly extracted features in the amplitude or frequency domain. Although effective to some extent, these approaches often fail to capture the dynamic interactions between brain regions. To overcome this limitation, brain connectivity analysis has gained attention, focusing on the relationships among different brain regions [horwitz2003elusive], [jang2018eeg], [jang2021eeg]. Connectivity matrices, constructed using metrics such as Pearson’s correlation coefficient (PCC), phase locking value (PLV) and transfer entropy (TE), provide a two-dimensional representation of interregional interactions and are widely used in emotional EEG classification [moon2020emotional].
@@ II-A EEG-Based Emotion Recognition
EEG has been widely utilized in emotion recognition due to its ability to capture real-time brain activity [moon2016implicit]. Both time-domain and frequency-domain analyses have been employed to interpret EEG signals associated with emotional states.
In particular, frequency-domain features have shown strong relevance to emotion classification. Liu et al. [liu2017real] distinguished emotional categories such as joy, anger, fear, and sadness using power spectral density (PSD) and asymmetry indices. Zhang et al. [zhang2016approach] further enhanced classification performance by incorporating time-frequency distribution analysis, which captures both temporal and spectral characteristics of EEG signals.
More recently, advances in deep learning have significantly improved feature extraction from EEG. Among various architectures, CNNs have demonstrated high performance in automated emotion recognition. Yanagimoto and Sugimoto [yanagimoto2016convolutional] applied CNNs to classify emotional valence and arousal levels, highlighting the effectiveness of learned representations over hand-crafted features. Li et al. [li2018hierarchical] proposed transforming EEG scalp maps into 2D images for CNN processing, enabling better handling of spatial and individual variability in emotional responses.
@@ II-B Brain Connectivity in Emotion Analysis
Costa et al. [costa2006eeg] demonstrated that phase synchronization patterns in EEG are significantly correlated with emotional states. Lee and Hsieh [lee2014classifying] proposed a method that integrates correlation, coherence, and phase synchronization to enhance emotion classification accuracy.
The importance of connectivity-based analysis is further supported by recent neuroimaging studies. Saarimäki et al. [saarimaki2022classification] showed that functional connectivity patterns from fMRI data reliably differentiate discrete emotion categories, especially within the default mode network. In the context of EEG, a recent systematic review by Khaleghi et al. [Khaleghi2024EEG] comprehensively analyzed how various brain abnormalities alter functional connectivity patterns, providing a broader clinical perspective on network disruptions.
Building on these findings, recent works have explored the integration of brain connectivity and deep learning. Moon et al. [moon2020emotional] demonstrated that EEG-derived connectivity matrices can serve as effective inputs to CNNs for learning emotion-related features, underscoring the potential of combining neural connectivity representations with data-driven models.
@@ III-A Connectivity Matrix Construction
$\text{PCC}(i,k)=\frac{\frac{1}{T}\sum_{t=1}^{T}(x_{i}^{t}-\mu_{i})(x_{k}^{t}-\mu_{k})}{\sigma_{i}\sigma_{k}}$ (1)
@@
Electrode placement within the connectivity matrix follows the distance-restricted arrangement method proposed in [moon2020emotional], which emphasizes spatial proximity among electrodes to reduce noise from long-range, spurious connections. In other words, electrodes are ordered such that those located in the same hemisphere and in close spatial proximity are positioned adjacently in the matrix. The resulting matrices are processed by CNNs to extract discriminative features for emotion classification.
@@ III-B Limitations of Conventional Approaches
This limitation arises primarily from the intrinsic nature of connectivity matrices, which encode pairwise relationships rather than spatially continuous features. For comparison, Fig. 2 presents a natural image from the CIFAR-10 dataset [krizhevsky2009learning] and its corresponding activation after the first convolutional layer of a CNN trained for object recognition. In this case, the model successfully captures meaningful structures such as edges and textures, as the input contains coherent spatial information. However, in the case of connectivity matrices, the absence of such spatial continuity impedes the CNN’s ability to learn effective representations.
@@ III-C Proposed Method
rr” refers to the upscaling factor.
@@ IV Experimental Setup
The evaluation in this study utilizes the DEAP dataset [koelstra2011deap], which contains EEG signals recorded from 32 participants while they viewed 40 emotional music videos. The EEG signals were initially recorded at a sampling rate of 512 Hz and subsequently downsampled to 128 Hz. The dataset includes emotion labels categorized by valence and arousal scales for each video stimulus. Based on the evaluation framework in [moon2020emotional], we develop a classification model to identify which of the 40 videos was being viewed based on the recorded EEG signals.
@@
As in [moon2020emotional], the one-minute EEG signal for each trial is divided into three-second-long segments. A simple non-overlapping segmentation would yield only 20 segments per trial. To augment the dataset, we apply an overlapping segmentation approach with a step size of 0.5 seconds, thereby increasing the number of samples. For each three-second segment, a 32×32×10 connectivity matrix is computed, capturing inter-channel relationships across the ten sub-frequency bands.
@@ V-B Visualization of t-SNE
To assess the discriminative quality of features extracted by the trained CNNs, we apply t-distributed stochastic neighbor embedding (t-SNE) [van2008visualizing] to the final feature vectors produced using different convolutional kernel sizes ($k=3,5,7$), under a fixed upscaling factor of ×3.0 with nearest-neighbor interpolation. The resulting embeddings are illustrated in Fig. 6.
@@ V-C Visualization of CNN Activation
To further understand how the input resolution affects feature learning, we visualize the activation maps from the first convolutional layer along with their corresponding Grad-CAM [selvaraju2017grad] outputs, which highlight the regions contributing most to the model’s final prediction. As illustrated in Fig. 6, higher upscaling factors lead to more spatially coherent and sharply defined activation patterns. This observation suggests that increased input resolution allows the model to more effectively exploit the spatial structure inherent in the connectivity matrices.
@@ VI Conclusion
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00453301) and by Institute of Information & communications Technology Planning & Evaluation (IITP) under 6G Cloud Research and Education Open Hub (IITP-2025-RS-2024-00428780) grant funded by the Korea government (MSIT). This work was also supported by the Yonsei Signature Research Cluster Program of 2024 (2024-22-0161).